-Solve the equation 8*+3 = 42x-1

please show steps so i can follow along
Answer #1

\small \textup{if }a^{x}=a^{y}\Rightarrow x=y

\small 8^{x+3}=4^{2x-1}

\small 2^{3.(x+3)}=2^{2.(2x-1)}

\small 3(x+3)=2(2x-1)

\small 3x+9=4x-2

\small x = 11